# Autonomous Smart Traffic Management System Using Artificial Intelligence CNN and LSTM

### v3

@@ Abstract
In the modern world, the development of Artificial Intelligence (AI) has contributed to improvements in various areas, including automation, computer vision, fraud detection, and more. AI can be leveraged to enhance the efficiency of Autonomous Smart Traffic Management (ASTM) systems and reduce traffic congestion rates. This paper presents an Autonomous Smart Traffic Management (STM) system that uses AI to improve traffic flow rates. The system employs the YOLO V5 Convolutional Neural Network to detect vehicles in traffic management images. Additionally, it predicts the number of vehicles for the next 12 hours using a Recurrent Neural Network with Long Short-Term Memory (RNN-LSTM). The Smart Traffic Management Cycle Length Analysis manages the traffic cycle length based on these vehicle predictions, aided by AI. From the results of the RNN-LSTM model for predicting vehicle numbers over the next 12 hours, we observe that the model predicts traffic with a Mean Squared Error (MSE) of 4.521 vehicles and a Root Mean Squared Error (RMSE) of 2.232 vehicles. After simulating the STM system in the CARLA simulation environment, we found that the Traffic Management Congestion Flow Rate with ASTM (21 vehicles per minute) is 50% higher than the rate without STM (around 15 vehicles per minute). Additionally, the Traffic Management Vehicle Pass Delay with STM (5 seconds per vehicle) is 70% lower than without STM (around 12 seconds per vehicle). These results demonstrate that the STM system using AI can increase traffic flow by 50% and reduce vehicle pass delays by 70%.
@@ I-A Artificial Intelligence
In 1997, AI regained hype after IBM’s Deep Blue, a chess-playing computer program, defeated reigning world chess champion and grandmaster Garry Kasparov [5]. However, it was not until 2012 that AlexNet, a CNN architecture by Alex Krizhevsky and his team, won the 2012 ImageNet annual image recognition challenge by a huge margin, reigniting development in the AI field [4]. To this day, numerous developments have emerged in AI, attributed to the vast amount of data available to train models, tremendous increases in computing power since the 1990s, and the introduction of revolutionary AI architectures with significantly higher performance and usability.

### v4

@@ Abstract
In the modern world, the development of Artificial Intelligence (AI) has contributed to improvements in various areas, including automation, computer vision, fraud detection, and more. AI can be leveraged to enhance the efficiency of Autonomous Smart Traffic Management (ASTM) systems and reduce traffic congestion rates. This paper presents an Autonomous Smart Traffic Management (STM) system that uses AI to improve traffic flow rates. The system employs the YOLO V5 Convolutional Neural Network to detect vehicles in traffic management images. Additionally, it predicts the number of vehicles for the next 12 hours using a Recurrent Neural Network with Long Short-Term Memory (RNN-LSTM). The Smart Traffic Management Cycle Length Analysis manages the traffic cycle length based on these vehicle predictions, aided by AI. From the results of the RNN-LSTM model for predicting vehicle numbers over the next 12 hours, we observe that the model predicts traffic with a Mean Squared Error (MSE) of 4.521 vehicles and a Root Mean Squared Error (RMSE) of 2.232 vehicles. After simulating the STM system in the CARLA simulation environment, we found that the Traffic Management Congestion Flow Rate with ASTM (21 vehicles per minute) is 50% higher than the rate without STM (around 15 vehicles per minute). Additionally, the Traffic Management Vehicle Pass Delay with STM (5 seconds per vehicle) is 70% lower than without ASTM (around 12 seconds per vehicle). These results demonstrate that the ASTM system using AI can increase traffic flow by 50% and reduce vehicle pass delays by 70%.
Keywords : Smart Traffic Management System , Automation, Artificial Intelligence, Computer Vision, Recurrent Neural Network, Traffic Simulation
@@ I-A Artificial Intelligence
In 1997, AI regained hype after IBM’s Deep Blue, a chess-playing computer program, defeated reigning world chess champion and grandmaster Garry Kasparov [5]. However, it was not until 2012 that AlexNet, a Convolutional Neural Network (CNN) architecture by Alex Krizhevsky and his team, won the 2012 ImageNet annual image recognition challenge by a huge margin, reigniting development in the AI field [4]. To this day, numerous developments have emerged in AI, attributed to the vast amount of data available to train models, tremendous increases in computing power since the 1990s, and the introduction of revolutionary AI architectures with significantly higher performance and usability.
@@ III Conclusion
List of Abbreviation
  • ASTM  : Autonomous Smart Traffic Management
  • AI  : Artificial Intelligence
  • ANN  : Artificial Neural Network
  • CNN  : Convolutional Neural Network
  • RNN  : Recurrent Neural Network
  • LSTM  : Long Short Term Memory
  • RMSE  : Root Mean Squared Error
@@ Acknowledgment
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data and experiment code and result will be made available on request.